# Dynamical stability of evolutionarily stable strategy in asymmetric games

### v1

@@ V Information theoretical aspects
The mention of KL-divergence (also, known as relative entropy) in the immediately preceding section is quite intriguing. The fact that a core concept of information theory is tied to the stability notion of ESS, reminds us of a known result in symmetric games in line with the general principle of minimum relative entropy Bomze (1991); Karev (2010); Qian et al. (1991); Floerchinger and Haas (2020): The KL-divergence of ESS from another time-evolving state monotonically decreases in symmetric games Baez and Pollard (2016). For completeness, let us recall that mathematically KL-divergence of a probability distribution with probability mass function (pmf) $p$ from another probability distribution ($q$ being the pmf) defined over the same random variable is expressed as
@@
What we have noted in the immediately preceding section is that $D_{\rm KL}(\mathbf{\hat{x}}||\mathbf{{x}})+D_{\rm KL}(\mathbf{\hat{y}}||\mathbf{{y}})$ decreases monotonically as the system races towards ESS of Definition 1a and 2a. However, what decreases in case of 2ESS is $D_{\rm KL}(\mathbf{\hat{x}}||\mathbf{{x}})+rD_{\rm KL}(\mathbf{\hat{y}}||\mathbf{{y}})$ (for some $r>0$). In passing, we may note that an even stricter condition on KL-divergence is compatible with Definition 1a: It is obvious that the fixed point corresponding to ESS of Definition 1a, leads to monotonic decrease of $D_{\rm KL}(\mathbf{\hat{x}}||\mathbf{{x}})$ and $D_{\rm KL}(\mathbf{\hat{y}}||\mathbf{{y}})$ separately.
@@
In the next step, we construct a Lyapunov function $V$ from the potential $U$. Let $V(\mathbf{x},\mathbf{y})=U(\mathbf{\hat{x}},\mathbf{\hat{y}})-U(\mathbf{x},\mathbf{y})$, if $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is the point where $U$ takes the maximum value, then $V(\mathbf{x},\mathbf{y})>0~{}\forall(\mathbf{x},\mathbf{y})\neq(\mathbf{\hat{x}},\mathbf{\hat{y}})$ at which $V$ is minimum—in fact, 00. Now, extrema of $V$ is attained at a point where
@@
This implies that $\dot{V}<0$ except at $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ where $\dot{V}=0$. Consequently, $V$ serves as the Lyapunov function indicating that $\mathbf{\hat{x}},\mathbf{\hat{y}}$ is locally asymptotically stable.
@@ VI ESS in Multiplayer hypermatrix games
In multiplayer games, the payoff hypermatrix depends on multiple arguments. We denote ${\sf U}[\mathbf{p}_{1};\mathbf{p}_{2},\cdots,\mathbf{p}_{\nu}]$ as the payoff of a player with strategy $\mathbf{p}_{1}$ interacting with other $\nu-1$ players with different strategies. If $d_{i}$ opponents are of same strategy, $p_{i}$ ($1<i\leq\nu$), we compactly denote the payoff as ${\sf U}[\mathbf{p}_{1};\mathbf{p}_{2}^{d_{2}},\cdots,\mathbf{p}_{\nu}^{d_{\nu}}]$; furthermore, if $d_{i}=1$, we suppress the superscript and $d_{i}=0$ means the absence of $p_{i}$-strategied opponent.
@@ VI.1 Asymmetric hypermatrix games
  1. Definition 1e: A state ($\mathbf{\hat{x}},\mathbf{\hat{y}}$) is an ESS in asymmetric hypermatrix game if for every population state $(\mathbf{x},\mathbf{y})$ that are sufficiently close but not equal to $(\mathbf{\hat{x}},\mathbf{\hat{y}})$,
    ${\sf A}[\mathbf{\hat{x}};\mathbf{x}^{\nu_{x}-1}]+{\sf B}[\mathbf{\hat{x}};\mathbf{y}^{\gamma_{x}}]>{\sf A}[\mathbf{x};\mathbf{x}^{\nu_{x}-1}]+{\sf B}\left[\mathbf{x};\mathbf{y}^{\gamma_{x}}\right]~{}\text{and}~{}{\sf C}[\mathbf{\hat{y}};\mathbf{x}^{\nu_{y}}]+{\sf D}[\mathbf{\hat{y}};\mathbf{y}^{\gamma_{y}-1}]>{\sf C}\left[\mathbf{y};\mathbf{x}^{\nu_{y}}\right]+{\sf D}[\mathbf{y};\mathbf{y}^{\gamma_{y}-1}].$ (67)
  2. Definition 2e: A state ($\mathbf{\hat{x}},\mathbf{\hat{y}}$) is an ESS in asymmetric hypermatrix game if for every population state $(\mathbf{x},\mathbf{y})$ that are sufficiently close but not equal to $(\mathbf{\hat{x}},\mathbf{\hat{y}})$,
@@
  3. Definition 3e: A state ($\mathbf{\hat{x}},\mathbf{\hat{y}}$) is an ESS in asymmetric hypermatrix game if for every population state $(\mathbf{x},\mathbf{y})$ that are sufficiently close but not equal to $(\mathbf{\hat{x}},\mathbf{\hat{y}})$,
    ${\sf A}[\mathbf{\hat{x}};\mathbf{x}^{\nu_{x}-1}]+{\sf B}[\mathbf{\hat{x}};\mathbf{y}^{\gamma_{x}}]>{\sf A}[\mathbf{x};\mathbf{x}^{\nu_{x}-1}]+{\sf B}\left[\mathbf{x};\mathbf{y}^{\gamma_{x}}\right]~{}\text{or}~{}{\sf C}[\mathbf{\hat{y}};\mathbf{x}^{\nu_{y}}]+{\sf D}[\mathbf{\hat{y}};\mathbf{y}^{\gamma_{y}-1}]>{\sf C}\left[\mathbf{y};\mathbf{x}^{\nu_{y}}\right]+{\sf D}[\mathbf{y};\mathbf{y}^{\gamma_{y}-1}].$ (69)
@@ VI.1.1 Pure ESS
ESS under Definition 1e for bihypermatrix games (${\sf A}={\sf D}={\sf 0}$) can only be pure. To prove this assertion, we recast the first inequality in (67) as
@@
This inequality has to hold for all values of $(\mathbf{x},\mathbf{y})$ in some neighbourhood of the ESS, $(\mathbf{\hat{x}},\mathbf{\hat{y}})$. If the ESS is not pure, then we can choose $x_{2}=\hat{x}_{2}$, $\cdots$, $x_{n-1}=\hat{x}_{n-1}$ and $y_{2}=\hat{y}_{2}$, $\cdots$, $y_{n-1}=\hat{y}_{n-1}$. Thus, one can always set the unfixed components $x_{1}$ and $y_{1}$ in such a manner that sign of the remaining term can be flipped; this, however, is not possible if the ESS is pure as $\hat{x}_{1}$ and $\hat{y}_{1}$ can be either zero or one. This proves the assertion. In fact, same assertion is valid for any ESS under Definition 2e (see Appendix B).
Of course, in general asymmetric hypermatrix games, ESS’s under all three definitions can be either pure or mixed. The next logical question is to ask ESS (of either bihypermatrix game or asymmetric hypermatrix game) under which definition is (locally) asymptotically stable fixed point of the replicator equation (66). As far as the pure ESS’s under Definitions 1e and 2e are concerned, the same Lyapunov function as used in the case of two-player game (see Sec. IV.1), $V_{1}(\mathbf{x},\mathbf{y})=D_{\rm KL}(\mathbf{\hat{x}}||\mathbf{{x}})+D_{\rm KL}(\mathbf{\hat{y}}||\mathbf{{y}})$, exists for the corresponding fixed points of Eq. (66). Thus, one can show that Definition 1e and Definition 2e imply the asymptotic stability of the (isolated) fixed point corresponding to a pure ESS—all one has to do is take the time derivative of $V_{1}$ and use Eq. (66) to arrive at
@@
and subsequently demand the $\dot{V}_{1}<0$ to validate the implication. Moreover, following exactly similar arguments as done in the two-player case (Sec. IV.1), linear stability analysis establishes that pure 2ESS need not be asymptotically stable fixed point of the replicator equation (66) even in the multiplayer case. We omit the straightforward steps as they are analogous and repetitive.
@@ VI.1.2 Mixed ESS
The above arguments using Lyapunov function, $V_{1}$, carry over to the mixed ESS’s under Definitions 1e and 2e, leading to the conclusion: Mixed ESS (except mixed 2ESS) is always locally asymptotically fixed point of replicator equation. The case of 2ESS is interesting.
First, we recall that interior fixed point can only be a center or saddle node in bimatrix games (see Sec. III.2). We now show that— extending the argument presented in Ref. Hofbauer and Sigmund (1998) for two player setting—similar assertion holds true even for multiplayer bihypermatrix games, i.e., an interior fixed point can neither be a source or sink. It is straightforward to see that, by definition, the internal fixed point ($\mathbf{\hat{x}},\mathbf{\hat{y}}$) of Eq. (66) with ${\sf A}={\sf D}=0$ satisfies the following condition:
@@
[…]
@@
In order to render calculations analytically tractable, we adopt the assumption that intraspecific interactions are infinitesimal, due to a small parameter $\mu$ that is multiplied to ${\sf A}$ and ${\sf D}$. Using the payoff hypermatrices of Fig. 2, we can rewrite Eq. (66) as
@@
Here, prime on a function denotes derivative with respective to the argument of that very function. Rather explicitly, we have
@@
Since signs of $\delta_{x}$ and $\delta_{y}$ can be positive or negative, at least one of the R.H.S.’s of Eq. (81) is positive in the entire neighbourhood; thus, the criteria of Definition 3e is satisfied but it does not meet the criteria of ESS of Definitions 1e or 2e. Thus, the locally asymptotically stable fixed point given by Eq. (79) is 2ESS. But does it satisfy inequality (76)? As we see below, it does not (when $\mu$ is very small as assumed in the calculations above).
Since we forced to do numerical check, we no longer need to assume $\mu$ to be infinitesimal. In other words, we now consider Eq. (77) with positive real $\mu$ and find its interior fixed points numerically for various values of $\mu$. We also validate that all the fixed points are 2ESS as well. For each such fixed points, we check if inequality (76) is satisfied. To investigate that, we divide the entire $(x,y)$ space into a uniform grid of size $50\times 50$. We then determine the sign of minimum value of L.H.S. of inequality (76)—denoted as $\min_{\forall(x,y)}(-\dot{V_{r}})$—for different values of $\mu$ and $r$ in entire $(x,y)$ grid. Positive $\min_{\forall(x,y)}(-\dot{V_{r}})$ implies that inequality (76) is satisfied. As seen in Fig. 3, we find a (green) region (in $r$-$\mu$ space discretized as $5000\times 1000$ grid in numerics) bounded by black solid curves, where $\min_{\forall(x,y)}(-\dot{V_{r}})>0$; elsewhere, we find that $\min_{\forall(x,y)}(-\dot{V_{r}})<0$ (see the red region in Fig. 3). At each fixed point, we also do linear stability analysis and find that all the fixed points under question are locally asymptotically stable. In conclusion, while 2ESS happens to be locally asymptotically stable, it need not satisfy Proposition 1 (inequality 76). We observe that if $\mu$ is very small, there is no such $r>0$ such that the inequality holds true for all $({x},{y})$; however, for higher values of $\mu$ such an $r$ exists.

### v2

@@ V Information theoretical aspects
The mention of KL-divergence (also, known as relative entropy) in the preceding sections is quite intriguing. The fact that a core concept of information theory is tied to the stability notion of ESS, reminds us of a known result in symmetric games in line with the general principle of minimum relative entropy Bomze (1991); Karev (2010); Qian et al. (1991); Floerchinger and Haas (2020): The KL-divergence of ESS from another time-evolving state monotonically decreases in symmetric games Baez and Pollard (2016). In asymmetric games, however, the presence of multiple definitions of ESS (the and-, sum-, and or-definitions) necessitates different adaptations of the KL-divergence to establish the relationship between KL-divergence and ESS. Furthermore, since Theorem 2 applies only to mixed states, the information-theoretic connection of the or-definition is also valid solely for mixed states.
For completeness, let us recall that mathematically KL-divergence of a probability distribution with probability mass function (pmf) $p$ from another probability distribution ($q$ being the pmf) defined over the same random variable is expressed as
@@
What we have noted in the immediately preceding section is that $D_{\rm KL}(\mathbf{\hat{x}}||\mathbf{{x}})+D_{\rm KL}(\mathbf{\hat{y}}||\mathbf{{y}})$ decreases monotonically as the system races towards ESS of and- and sum-definitions. However, what decreases in case of mixed 2ESS( or-definition) is $D_{\rm KL}(\mathbf{\hat{x}}||\mathbf{{x}})+rD_{\rm KL}(\mathbf{\hat{y}}||\mathbf{{y}})$ (for some $r>0$). In passing, we may note that an even stricter condition on KL-divergence is compatible with and-definition: It is obvious that the fixed point corresponding to ESS of and-definition, leads to monotonic decrease of $D_{\rm KL}(\mathbf{\hat{x}}||\mathbf{{x}})$ and $D_{\rm KL}(\mathbf{\hat{y}}||\mathbf{{y}})$ separately.
@@
In the next step, we construct a Lyapunov function $V$ from the potential $U$. Let $V(\mathbf{x},\mathbf{y})=U(\mathbf{\hat{x}},\mathbf{\hat{y}})-U(\mathbf{x},\mathbf{y})$, if $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is the point where $U$ takes the maximum value, then $V(\mathbf{x},\mathbf{y})>0~{}\forall(\mathbf{x},\mathbf{y})\neq(\mathbf{\hat{x}},\mathbf{\hat{y}})$ and at $(\mathbf{\hat{x}},\mathbf{\hat{y}})$, $V$ takes the minimum value—in fact, 00. Now, extrema of $V$ is attained at a point where
@@
This implies that $\dot{V}<0$ except at $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ where $\dot{V}=0$. Consequently, $V$ serves as the Lyapunov function indicating that $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is locally asymptotically stable.
@@ VI ESS in Multiplayer hypermatrix games
In multiplayer games, the payoff hypermatrix depends on multiple arguments. We denote ${\sf U}[\mathbf{p}_{1};\mathbf{p}_{2},\cdots,\mathbf{p}_{\nu}]$ as the payoff of a player with strategy $\mathbf{p}_{1}$ interacting with other $\nu-1$ players with different strategies. If $d_{i}$ opponents are of same strategy, $\mathbf{p}_{i}$ ($1<i\leq\nu$), we compactly denote the payoff as ${\sf U}[\mathbf{p}_{1};\mathbf{p}_{2}^{d_{2}},\cdots,\mathbf{p}_{\nu}^{d_{\nu}}]$; furthermore, if $d_{i}=1$, we suppress the superscript and $d_{i}=0$ means the absence of $\mathbf{p}_{i}$-strategied opponent.
@@ VI.1 Asymmetric hypermatrix games
  1. Definition 1e (and-definition): A state ($\mathbf{\hat{x}},\mathbf{\hat{y}}$) is an ESS in asymmetric hypermatrix game if for every population state $(\mathbf{x},\mathbf{y})$ that are sufficiently close but not equal to $(\mathbf{\hat{x}},\mathbf{\hat{y}})$,
    ${\sf A}[\mathbf{\hat{x}};\mathbf{x}^{\nu_{x}-1}]+{\sf B}[\mathbf{\hat{x}};\mathbf{y}^{\gamma_{x}}]>{\sf A}[\mathbf{x};\mathbf{x}^{\nu_{x}-1}]+{\sf B}\left[\mathbf{x};\mathbf{y}^{\gamma_{x}}\right]~{}\text{\emph{and}}~{}{\sf C}[\mathbf{\hat{y}};\mathbf{x}^{\nu_{y}}]+{\sf D}[\mathbf{\hat{y}};\mathbf{y}^{\gamma_{y}-1}]>{\sf C}\left[\mathbf{y};\mathbf{x}^{\nu_{y}}\right]+{\sf D}[\mathbf{y};\mathbf{y}^{\gamma_{y}-1}].$ (72)
  2. Definition 2e (sum-definition): A state ($\mathbf{\hat{x}},\mathbf{\hat{y}}$) is an ESS in asymmetric hypermatrix game if for every population state $(\mathbf{x},\mathbf{y})$ that are sufficiently close but not equal to $(\mathbf{\hat{x}},\mathbf{\hat{y}})$,
@@
  3. Definition 3e (or-definition): A state ($\mathbf{\hat{x}},\mathbf{\hat{y}}$) is an ESS in asymmetric hypermatrix game if for every population state $(\mathbf{x},\mathbf{y})$ that are sufficiently close but not equal to $(\mathbf{\hat{x}},\mathbf{\hat{y}})$,
    ${\sf A}[\mathbf{\hat{x}};\mathbf{x}^{\nu_{x}-1}]+{\sf B}[\mathbf{\hat{x}};\mathbf{y}^{\gamma_{x}}]>{\sf A}[\mathbf{x};\mathbf{x}^{\nu_{x}-1}]+{\sf B}\left[\mathbf{x};\mathbf{y}^{\gamma_{x}}\right]~{}\text{\emph{or}}~{}{\sf C}[\mathbf{\hat{y}};\mathbf{x}^{\nu_{y}}]+{\sf D}[\mathbf{\hat{y}};\mathbf{y}^{\gamma_{y}-1}]>{\sf C}\left[\mathbf{y};\mathbf{x}^{\nu_{y}}\right]+{\sf D}[\mathbf{y};\mathbf{y}^{\gamma_{y}-1}].$ (74)
@@ VI.1.1 Pure ESS
ESS under and-definition for bihypermatrix games (${\sf A}={\sf D}={\sf 0}$) can only be pure. To prove this assertion, we recast the first inequality in (72) as
@@
This inequality has to hold for all values of $(\mathbf{x},\mathbf{y})$ in some neighbourhood of the ESS, $(\mathbf{\hat{x}},\mathbf{\hat{y}})$. If the ESS is not pure, then we can choose $x_{2}=\hat{x}_{2}$, $\cdots$, $x_{n-1}=\hat{x}_{n-1}$ and $y_{2}=\hat{y}_{2}$, $\cdots$, $y_{n-1}=\hat{y}_{n-1}$. Thus, one can always set the unfixed components $x_{1}$ and $y_{1}$ in such a manner that sign of the remaining term can be flipped; this, however, is not possible if the ESS is pure as $\hat{x}_{1}$ and $\hat{y}_{1}$ can be either zero or one. This proves the assertion. In fact, same assertion is valid for any ESS under sum-definition (see Appendix B).
Of course, in general asymmetric hypermatrix games, ESS’s under all three definitions can be either pure or mixed. The next logical question is to ask ESS (of either bihypermatrix game or asymmetric hypermatrix game) under which definition is (locally) asymptotically stable fixed point of the replicator equation (71). As far as the pure ESS’s under and- and sum-definitions are concerned, the same Lyapunov function as used in the case of two-player game (see Sec. III.2), $V_{1}(\mathbf{x},\mathbf{y})=D_{\rm KL}(\mathbf{\hat{x}}||\mathbf{{x}})+D_{\rm KL}(\mathbf{\hat{y}}||\mathbf{{y}})$, exists for the corresponding fixed points of Eq. (71). Thus, one can show that and- and sum-definitions imply the asymptotic stability of the (isolated) fixed point corresponding to a pure ESS—all one has to do is take the time derivative of $V_{1}$ and use Eq. (71) to arrive at
@@
and subsequently demand the $\dot{V}_{1}<0$ to validate the implication. Moreover, following exactly similar arguments as done in the two-player case (Sec. IV.1), linear stability analysis establishes that pure 2ESS is asymptotically stable fixed point of the replicator equation (71) even in the multiplayer case when it is a strict NE. Using the same argument we can also show that all three definitions are equivalent in case of strict NE. We omit the straightforward steps as they are analogous and repetitive, as one can see from the eigenvalues $\lambda_{1},\lambda_{2},\cdots,\lambda_{m-1},\lambda^{\prime}_{1},\lambda^{\prime}_{2},\cdots,\lambda^{\prime}_{n-1}$—which respectively are— $({\sf A}[\mathbf{p}_{1};\mathbf{p}_{m}^{\nu_{x}-1}]+{\sf B}[\mathbf{p}_{1};\mathbf{q}_{n}^{\gamma_{x}}]-{\sf A}[\mathbf{p}_{m};\mathbf{p}_{m}^{\nu_{x}-1}]-{\sf B}[\mathbf{p}_{m};\mathbf{q}_{n}^{\gamma_{x}}]),({\sf A}[\mathbf{p}_{2};\mathbf{p}_{m}^{\nu_{x}-1}]+{\sf B}[\mathbf{p}_{2};\mathbf{q}_{n}^{\gamma_{x}}]-{\sf A}[\mathbf{p}_{m};\mathbf{p}_{m}^{\nu_{x}-1}]-{\sf B}[\mathbf{p}_{m};\mathbf{q}_{n}^{\gamma_{x}}]),\cdots,({\sf A}[\mathbf{p}_{m-1};\mathbf{p}_{m}^{\nu_{x}-1}]+{\sf B}[\mathbf{p}_{m-1};\mathbf{q}_{n}^{\gamma_{x}}]-{\sf A}[\mathbf{p}_{m};\mathbf{p}_{m}^{\nu_{x}-1}]-{\sf B}[\mathbf{p}_{m};\mathbf{q}_{n}^{\gamma_{x}}]),({\sf C}[\mathbf{q}_{1};\mathbf{p}_{m}^{\nu_{y}}]+{\sf D}[\mathbf{q}_{1};\mathbf{q}_{n}^{{\gamma_{y}}-1}]-{\sf C}[\mathbf{q}_{n};\mathbf{p}_{m}^{\nu_{y}}]-{\sf D}[\mathbf{q}_{n};\mathbf{q}_{n}^{\gamma_{y}-1}]),({\sf C}[\mathbf{q}_{2};\mathbf{p}_{m}^{\nu_{y}}]+{\sf D}[\mathbf{q}_{2};\mathbf{q}_{n}^{\gamma_{y}-1}]-{\sf C}[\mathbf{q}_{n};\mathbf{p}_{m}^{\nu_{y}}]-{\sf D}[\mathbf{q}_{n};\mathbf{q}_{n}^{\gamma_{y}-1}]),\cdots,({\sf C}[\mathbf{q}_{n-1};\mathbf{p}_{m}^{\nu_{y}}]+{\sf D}[\mathbf{q}_{n-1};\mathbf{q}_{n}^{\gamma_{y}-1}]-{\sf C}[\mathbf{q}_{n};\mathbf{p}_{m}^{\nu_{y}}]-{\sf D}[\mathbf{q}_{n};\mathbf{q}_{n}^{\gamma_{y}-1}])$, at the fixed point $(\hat{\mathbf{x}},\hat{\mathbf{y}})=((0,\cdots,0,1),(0,\cdots,0,1))$.
It is not very hard to see that, for ESS that is non-strict pure NE, the situation is more nuanced than for ESS that is strict NE. In fact, similar or even more intricate subtleties may arise, as observed in two-player games.
@@ VI.1.2 Mixed ESS
The above arguments using Lyapunov function, $V_{1}$, carry over to the mixed ESS’s under and- and sum-definitions as well, leading to the conclusion: Mixed ESS (except mixed 2ESS) is always locally asymptotically fixed point of replicator equation. The case of 2ESS (or-definition) is interesting.
First, we recall that interior fixed point can only be a center or saddle node in bimatrix games (see Sec. III.3). We now show that— extending the argument presented in Ref. Hofbauer and Sigmund (1998) for two player setting—similar assertion holds true even for multiplayer bihypermatrix games, i.e., an interior fixed point can neither be a source or sink. It is straightforward to see that, by definition, the internal fixed point ($\mathbf{\hat{x}},\mathbf{\hat{y}}$) of Eq. (71) with ${\sf A}={\sf D}=0$ satisfies the following condition:
@@
Next, we come to mixed 2ESS in asymmetric hypermatrix game. One could expect that as done previously in Sec. IV.2, $V_{r}(\mathbf{x},\mathbf{y})=D_{\rm KL}(\mathbf{\hat{x}}||\mathbf{{x}})+rD_{\rm KL}(\mathbf{\hat{y}}||\mathbf{{y}})$ as a candidate for Lyapunov function. For this function to be a valid Lyapunov function one needs an analog of Theorem 2 for multiplayer games, that is, the following proposition needs to be true:
@@
In order to render calculations analytically tractable, we adopt the assumption that intra-specific interactions are infinitesimal, due to a small parameter $\mu$ that is multiplied to ${\sf A}$ and ${\sf D}$. Using the payoff hypermatrices of Fig. 2, we can rewrite Eq. (71) as
@@
Here, prime on a function denotes derivative with respect to the argument of that very function. Rather explicitly, we have
@@
Since signs of $\delta_{x}$ and $\delta_{y}$ can be positive or negative, at least one of the R.H.S.’s of Eq. (86) is positive in the entire neighbourhood; thus, the criteria of or-definition is satisfied but it does not meet the criteria of ESS of and- and sum-definitions. Thus, the locally asymptotically stable fixed point given by Eq. (84) is 2ESS. But does it satisfy inequality (81)? As we see below, it does not (when $\mu$ is very small as assumed in the calculations above).
Since we forced to do numerical check, we no longer need to assume $\mu$ to be infinitesimal. In other words, we now consider Eq. (82) with positive real $\mu$ and find its interior fixed points numerically for various values of $\mu$. We also validate that all the fixed points are 2ESS as well. For each such fixed points, we check if inequality (81) is satisfied. To investigate that, we divide the entire $(x,y)$ space into a uniform grid of size $50\times 50$. We then determine the sign of minimum value of L.H.S. of inequality (81)—denoted as $\min_{\forall(x,y)}(-\dot{V_{r}})$—for different values of $\mu$ and $r$ in entire $(x,y)$ grid. Positive $\min_{\forall(x,y)}(-\dot{V_{r}})$ implies that inequality (81) is satisfied. As seen in Fig. 3, we find a (green) region (in $r$-$\mu$ space discretized as $5000\times 1000$ grid in numerics) bounded by black solid curves, where $\min_{\forall(x,y)}(-\dot{V_{r}})>0$; elsewhere, we find that $\min_{\forall(x,y)}(-\dot{V_{r}})<0$ (see the red region in Fig. 3). At each fixed point, we also do linear stability analysis and find that all the fixed points under question are locally asymptotically stable. In conclusion, while 2ESS happens to be locally asymptotically stable, it need not satisfy Proposition 1 [inequality (81)]. We observe that if $\mu$ is very small, there is no such $r>0$ such that the inequality holds true for all $({x},{y})$; however, for higher values of $\mu$ such an $r$ exists.